OpenStudy (callisto):

Level 3 question - locus again

OpenStudy (callisto):

lol, I forgot to delete something :S

OpenStudy (anonymous):

Hmm Co-ordinate of Q be (b,0) and P (b,a). Equation of L is 3x + 4y + 14= 0. So, 3b + 4a + 14 = 0 (1). You have the relation for a and b now. \[N\equiv \left(b,\frac{2a}5\right)\] \[x = b\tag2\]\[5y = 2a\tag3\] Now solve the equations 1,2 and 3.

OpenStudy (anonymous):

Is Chinese Cartesian system different?

OpenStudy (thomas5267):

Why you asked that question?

OpenStudy (anonymous):

While I defined Q as (b,0). She defined it as (0.b). http://assets.openstudy.com/updates/attachments/4f8be056e4b0935d57e9b307-callisto-1334567021212-math.jpg I should say co-ordinate system, right? Because the one we follow is Cartesian co-ordinate system.

OpenStudy (callisto):

Q is (0,b) ...

OpenStudy (anonymous):

*facepalm* :(

OpenStudy (anonymous):

*epic facepalm* My co-ordinate points aren't right but the concept is right you should use the concept.

OpenStudy (anonymous):

Or, N is now (2a/5,b) and the relation between a and b is 3a + 4b = 14. Maybe.

OpenStudy (callisto):

BTW, that wasn't my workings...

OpenStudy (callisto):

Can I do it in this way? Let N=(x,y) , Q = (0,b) Since P is perpendicular to Q, y coordinate of P =b Put y=b into 3x+4y+14 =0 3x + 4b +14 =0 x = (-4b-14)/3 So, P = [ (-4b-14)/3 , b] By section formula [(-4b-14)/3 ] x3 + 0 x2 ------------------- =x 5 (-4b-14) =5x b = (5x+14)/(-4) b x3 + b x2 ----------- =y 5 y = b = (5x+14)/(-4) Therefore, the equation is y =(5x+14)/(-4)

OpenStudy (anonymous):

i'm doing it differently but yours looks good...

OpenStudy (callisto):

What's the way you're using?

OpenStudy (anonymous):

i'm using distance like the ellipse locus earlier...

OpenStudy (callisto):

Basically, I just 'copied' the way I did for the previous locus problem

OpenStudy (callisto):

lol

OpenStudy (anonymous):

and that's what i'm doing with my previous one too..

OpenStudy (anonymous):

***this one

OpenStudy (callisto):

This time, you'll probably give an elegant solution like last time :)

OpenStudy (anonymous):

dunno... my "y" keeps vanishing!

OpenStudy (anonymous):

3a + 4b = - 14, N (x,y) \(\equiv \left(\frac{2a}5, b\right)\) \[3 a\frac22 \cdot \frac55 + 4\cdot b = 14\implies \frac52 \cdot 3x + 4y +14=0 \]Is this wrong?

OpenStudy (callisto):

/_\ .... @Ishaan94 you know you're always better than me.... and I dare not challenge you....:(

OpenStudy (anonymous):

So the answer is not matching, let me check it again. I commit lots of silly mistakes. :/

OpenStudy (anonymous):

read this attach file..

OpenStudy (anonymous):

Hmm Callisto was right and I wrong.